# JWST Observations of the Double Nucleus in NGC 4486B: Possible Evidence for a Recent Binary SMBH Merger and Recoil

### v1

@@ Abstract
A recent study of the compact elliptical galaxy NGC 4486B using JWST–NIRSpec IFU kinematics confirmed a supermassive black hole (SMBH) of mass ${\cal M}_{BH}=3.6\pm 0.7\times 10^{8}$ ($\sim 8$% of the stellar mass). In addition to its double nucleus, the nuclear kinematics show pronounced asymmetries: a velocity-dispersion peak displaced by 6 pc from the galaxy center and a $\sim 16\,\mathrm{km\ s}^{-1}$ offset in the mean stellar line-of-sight velocity near the SMBH. We examine the origin of the 12 pc double nucleus and these asymmetries and show that the observations favor an SMBH surrounded by an eccentric nuclear disk (END). END formation models require the SMBH to experience a GW recoil following a binary SMBH merger. Our orbit-superposition models contain $\sim 50\%$ retrograde stars at the edge of the nuclear region, in striking agreement with END-formation simulations. We infer a pre-merger mass ratio $q>0.15$ and a recoil kick of $\sim 340\,\mathrm{km\ s}^{-1}$. Our $N$-body simulations show that with such a kick, the SMBH returns to the center within $\sim 30$ Myr. Its flat central core is also consistent with earlier “binary black hole scouring”. We test two alternative mechanisms—buoyancy-driven oscillations and a pre-merger SMBH binary—but neither reproduces the observed offsets, favoring the GW-kick scenario. Our direct $N$-body simulations further show that a prograde SMBH binary in a rotating host can stall in a corotation resonance, delaying coalescence. Thus, although NGC 4486B is an old, relaxed galaxy near the Virgo cluster center, its SMBH appears to have merged only recently, making its nucleus a rare nearby laboratory for studying post-merger SMBH dynamics.
@@
1 Introduction
NGC 4486B (VCC1279) is a cE galaxy at the center of the Virgo Cluster, projected $7.3^{\prime}$ ($\sim 35$ kpc) from the massive elliptical M87, with a modest line-of-sight velocity difference of $\sim 200\,\mathrm{km}\,\mathrm{s^{-1}}$ that indicates a close dynamical association. NGC 4486B is characterized by a small effective radius of $2.33^{\prime\prime}$ ($\sim 0.19$ kpc) and an absolute magnitude of $M_{V}\sim 17.6$, with a total stellar mass of $_{*}\sim 6\times 10^{9}MM_{\odot}$ (Lauer.1996; Ferrarese.2006; Janz.2016). Recently, we showed using JWST-NIRSpec IFU kinematical data and archival HST photometry, that NGC 4486B hosts a supermassive black hole (SMBH) with a mass of ${\cal M}_{BH}=(3.6\pm 0.7)\times 10^{8}\mbox{$M_{\odot}$}$, making this black hole significantly overmassive (${\cal M}_{BH}/M_{*}\sim 8$%) relative to expectations from black hole scaling relations (Behzad.2025, hereafter Paper 1). This suggests that NGC 4486B is an extreme case of a tidally stripped galaxy, likely the compact core of a once significantly more massive system.
@@ 1 Introduction
One plausible explanation for the double nucleus in M31 and NGC4486B is the presence of an eccentric, apsidally aligned stellar disk orbiting a central SMBH (Tremaine.1995; Peiris.2003; Bacon.2001; Sambhus.2002; BrownMagorrian.2013; Wernke.2021). In this model, the nearly Keplerian potential of the SMBH dominates the stellar dynamics, shaping a lopsided configuration where stars follow coherent, eccentric orbits. The apparent double nucleus arises from the asymmetric distribution of stellar density along these orbits: the brighter component corresponds to stars near their apocenters, where they move more slowly, while the fainter component is associated with stars near their pericenters, close to the black hole.
@@
Lauer.1996 explored a range of other possible origins for the double nucleus in NGC 4486B. A dust absorption feature as well as an infalling globular cluster were ruled out due to the nearly identical morphologies in $V$ and $I$ bands and the lack of significant color gradients across the nucleus. An ongoing merging or binary stellar nucleus was also deemed unlikely, as the decay timescale for such a system due to dynamical friction is short ($\lesssim 10^{7}$ yr), and there is no evidence for a recent merger in NGC 4486B. Projection effects from unrelated foreground or background sources were likewise dismissed based on the nearly identical colors and recession velocities of the two nuclei (Kormendy.1997).
@@
A striking feature of NGC4486B is its flat central density core (see Section 2.1). One of the best accepted mechanisms for producing such cores in early type galaxies is via “binary-black-hole scouring” following the merger of two gas poor galaxies, each with its own SMBH. The black holes (BHs) quickly sink to the common center of the galaxy via dynamical friction and become bound with separations of 10s of parsecs. In order to shrink the orbit further, the SMBH binary transfers its orbital kinetic energy to surrounding stars, via the $3$-body slingshot effect, preferentially ejecting stars with low angular momentum (e.g. Milos.2003; Baile15) until the binary separation becomes small enough for gravitational wave (GW) radiation to dominate. High-resolution $N$-body simulations (khan+12a; Rantala2024) have shown that an SMBH binary can eject a stellar mass amounting to 252-5 times its total mass from the central region, thereby converting a central cusp of stars into a flatter central core (rantala2018; Nasim2021; Khan2025).
@@
Numerical relativity simulations (e.g., Campanelli.2007ApJ; Campanelli.2007PRL) have shown that the anisotropic emission of GWs during the merger of two SMBH delivers a kick to the final black hole with a magnitude that depends on a variety of factors: the mass ratio $q=m_{2}/M_{1}$ of the secondary to the primary; whether the BHs have spin; the alignment of their spins to their mutual orbital plane; and whether the black hole spins are aligned or misaligned. While kicks as large as $\sim 2000-4000\,\mathrm{km\ s}^{-1}$ (Campanelli.2007PRL) are possible, the actual distribution of kick velocities is uncertain and most kicks are probably $\lesssim 1000$$\mathrm{km\ s}^{-1}$(Gualandris.2008). Even a modest kick to the SMBH causes it to oscillate transfering orbital energy to nearby stars, further flattening the central density profile (Gualandris.2008; Khonji.2024).
@@
Recently, the eccentric nuclear disk (END) model has received renewed interest due to a suite of direct $N$-body simulations of nuclear disks surrounding SMBHs (Akiba.2021; Akiba.2023; Bright.2024). These authors demonstrate that a GW recoil kick can naturally transform a circular stellar disk bound to an SMBH into an eccentric, apsidally aligned lopsided one. In this scenario, the stars that remain bound to the SMBH after the kick experience angular-momentum changes that place them on eccentric orbits, whose net alignment is oriented roughly orthogonal to the kick direction. For a sufficiently large kick, stars in the outer parts of the disk can even become retrograde in the rest frame of the SMBH.
@@
In this work, we investigate the photometric and kinematic signatures of the double nucleus in NGC 4486B and build upon our previous dynamical modeling results from Paper 1, as well as previous $N$-body simulations, to examine the origin and dynamical state of its END. Using a self-consistent $N$-body realization constructed using our Schwarzschild orbit-superposition model from Paper 1, we perform simulations exploring both pre- and post-merger scenarios. We use previous analysis from Akiba.2021; Akiba.2023; Akiba.2025 and the observed properties of the double nucleus to estimate the kick velocity and likely mass-ratio of the black hole binary. We simulate various recoil kicks to assess the timescale on which the black hole would sink to the center. In addition to GW recoil, we test other alternative mechanisms capable of displacing the SMBH and influencing the observed kinematics, such as buoyancy-driven displacements of the SMBH (Cole2012; Banik2021; Banik2022; Dattathri2025a), and the pre-merger evolution of a prograde binary black hole merger (holley+15; Rasskazov+16; Mirza+17; Khan2020). We also quantify the expected amplitude of SMBH Brownian motion (Merritt.2001) to verify that the displacements seen in our simulations and inferred from observations exceed the stochastic wandering expected from finite stellar encounters Brownian motion.
@@
While nearly two decades of theoretical and simulation work have explored the consequences of binary black hole mergers, the resulting GW recoils, and SMBH Brownian motion in galaxies, NGC 4486B appears to be the first system exhibiting multiple observable signatures of a recent merger—making it a benchmark case for studying the dynamical aftermath of SMBH coalescence. Although NGC 4486B and M31 are among the best-studied examples, double nuclei are not unique. Debattista.2006 examined a sample of dE galaxies from Lotz.2004 imaged in the WFPC2 snapshot surveys GO-8600 and GO-6352 (PI: H. Ferguson) and found that VCC 1107, FCC 208, and VCC 128 show signs of double nuclei, with VCC 128 exhibiting two comparably bright components and the closest separation ($\sim 32$ pc). Tracing the formation and evolution of the eccentric disk in NGC 4486B and other galaxies may provide indirect constraints on the frequency of SMBH mergers – an essential parameter for interpreting the stochastic GW background measured by Pulsar Timing Arrays (PTAs) (Nanograv.2023; nanograv.2024; EPTA.2023; EPTA.2024).
@@
The structure of this paper is as follows. In Section 2, we describe the HST and JWST datasets used in our analysis and briefly summarize the Schwarzschild dynamical modeling from Paper 1. In Section 3, we present the observational evidence for an END in NGC 4486B and compare the nuclear structure and kinematics of our Schwarzschild model with simulations of ENDs formed through GW recoil kicks. Using the observed properties of the double nucleus, we derive constraints on the black hole binary mass ratio and the required recoil velocity. In Section 4, we use direct $N$-body simulations to follow the post-kick trajectory of the SMBH and evaluate the timescale on which it returns to the galaxy center, as well as the kinematic signatures such a kick would imprint. Section 5 examines alternative mechanisms that could produce an off-center SMBH or similar kinematic asymmetries. Finally, in Section 6 we summarize our findings and discuss their implications. Throughout this study, we adopt a distance of $16.3$ Mpc to NGC 4486B (Blakeslee.2009)
2 Observational data and overview of Schwarzschild models in Paper 1
@@
In this section we briefly summarize the properties of photometric imaging from Hubble Space Telescope (HST) instruments WFPC2 and ACS/WFC and the spectroscopic data from JWST-NIRSpec IFU that were used to construct the Schwarzschild dynamical models of NGC 4486B presented in Paper 1 (Behzad.2025). In addition, we briefly summarize the main results from that paper including details about the best-fit Schwarzschild models that are used in this work.
2.1 Photometric data and image analysis
@@
In Paper 1, we used archival HST-ACS imaging in the F850LP filter (pixel scale $\sim 0.05\arcsec$ $\lesssim 4$ pc and PSF FWHM $\sim 0.9\arcsec$ $\lesssim 72$ pc) acquired as part of the ACS Virgo Cluster Survey (Cote.2004). A model point spread function (PSF) was generated using the TinyTim software package (Krist.2011) and used for all photometric analyses. The full photometric frame reaches out to $\sim$1 kpc in radius, with an effective radius of $2.33\arcsec$ (182 pc), extending well beyond the field of view of the NIRSpec-IFU observations. We also examine the WFPC2 F [^5] W image from Lauer.1996, which provides slightly higher spatial resolution in the central region and reveals the double nucleus structure more distinctly.
Previous analyses of the morphology, isophotal parameters, and surface brightness profile of NGC 4486B showed that it exhibits a shallow surface brightness profile in its central region (Lauer.1996; Faber.1997; Ferrarese.2006). Ferrarese.2006 modeled its surface brightness using a single Sérsic component and obtained $n=2.73$ and $r_{e}=2.33\arcsec$. We repeated the 2D surface brightness fitting of the ACS F850LP image using GALFIT (GALFIT2010) models with one, two, three, and four Sérsic components. For a single component, the Sérsic index is $n=3.05$, slightly higher than in previous work. As the number of Sérsic components increases to three, the quality of the fit improved. Beyond this point, however, the model did not show substantial improvement (see Table 1). By adding more components, the Sésic index of the inner region decreases to a range of n=1.21.4n=1.2-1.4. As might be expected due to the double nuclei, the inner region of NGC 4486B cannot be adequately described by single or multiple nested Sérsic components (even allowing for the components to have different centers). The major-axis position angles of all components across different models are similar and consistent with the position angle profile derived from the ellipse fitting to the isophotes. The ellipticity and position angle of the isophotes (outside the innermost $\sim 0.15\arcsec$ region occupied by the double nucleus) remain relatively constant with radius, indicating that the galaxy is close to axisymmetric overall. Similar results were obtained with the WFPC2 F [^5] W image.
@@
Figure 2: First row, right panel: Unsharp-masked WFPC2/F [^5] W image of the double nucleus in NGC 4486B, showing two peaks separated by $\sim$12 pc. First row, left and middle panels: Stellar kinematic maps of $v$ and $\sigma$ from JWST-NIRSpec IFU data. In all panels, surface-brightness contours from the deconvolved WFPC2/F [^5] W image are overplotted. The black dot marks the center of the faint peak (P2), which coincides with the $\sigma$ peak from NIRSpec and is identified as the likely location of the SMBH. The purple dot indicates the center of the bright peak (P1), coinciding with the center of the IFU data cube. The vertical dashed line marks the kinematic center. Second row: One-dimensional kinematic profiles of NGC 4486B derived from JWST-NIRSpec IFU data, showing the $|v|$, $\sigma$, and $|v|/\sigma$ values along the major axis intersecting both nuclei. Red points represent measurements and associated uncertainties from the right side of the kinematic maps, where $\sigma$ peaks. Blue points correspond to the left side of the galaxy, mirrored about the rotation axis for direct comparison. The gray shaded region marks the approximate radial extent of the eccentric disk inferred from photometric and kinematic features (corresponds to the white rectangle above).
In Paper 1, we used the multi-Gaussian expansion (MGE) method (Cappellari.2002) to model the surface brightness of NGC 4486B. The details of the fit are provided in Table 1 of Paper 1. In the top panel of Figure 1, we present the projected luminosity density of NGC 4486B (black dots), along with the corresponding models: the MGE model (in red) and Sérsic models with one (green), two (blue), three (orange), and four (gray) components. The vertical dashed line indicates the PSF FWHM of the F850LP image. None of the models with multiple Sérsic components are able to replicate the flatness of the inner profile as well as the MGE model does in this region. To test whether the presence of the double nucleus affects the derived surface brightness profile, we allowed the centers of the Sérsic (or Gaussian) components to vary freely in our GALFIT modeling. In the best-fit model, the centers of all components are found to lie within $\leq 1$ pixel of each other. Likewise, in the standard MGE fitting, where all Gaussian components share a fixed center at the brightest peak, the result remains consistent. The choice of fixing or freeing the component centers therefore does not significantly affect the shape of the final luminosity profile. This is likely because, although the nuclear region is slightly lopsided rather than fully symmetric, the two peaks are separated by only $\simeq$ 3-4 pixels, comparable to the PSF FWHM. Consequently, the central light distribution appears smooth rather than distinctly double in the F850LP image, causing all model components to converge to nearly the same photocenter, which lies on the bright peak.
@@
As mentioned in Section 1, the favored mechanism for core formation in elliptical galaxies is “core scouring by an SMBH binary”. The low-density core produced in this manner can extend up to five times the black hole binary influence radius (rantala2018; Nasim2021; Khan2025). For the black hole in NGC 4486B ${\cal M}_{BH}=3.6\times 10^{8}$$M_{\odot}$ (Paper 1), the sphere of influence is estimated to be 203520-35 pc, which is comparable to the observed core radius.
The surface brightness fitting above ignored the double nucleus, which is barely resolved in the ACS/WFC F850LP image. In the higher spatial resolution of the HST WFC2 F [^5] W image (Lauer.1996) however, it is clearly resolved (Figure 1 of Paper 1 gives both images). Figure 2 (first row, right panel) shows an unsharp-masked image of NGC 4486B, in which the double nucleus is revealed with striking clarity. The two peaks are distinctly separated by $\sim$12 pc, with the fainter component (P2) marked by a black dot while the brighter nucleus (P1) is marked by a purple dot. Photometrically, P1 has a central surface brightness of $\mu_{I}=12.89$ mag arcsec-2, while P2 has $\mu_{I}=13.00$ mag arcsec-2, with a minimum of $\mu_{I}=13.06$ mag arcsec-2 between them (Lauer.1996). The white rectangle in this image encloses the nuclear region that we consider in greater detail below.
2.2 JWST NIRSpec-IFU kinematics

### v2

@@ Abstract
A recent study of the compact elliptical galaxy NGC 4486B using JWST–NIRSpec IFU kinematics confirmed a supermassive black hole (SMBH) of mass ${\cal M}_{BH}=3.6\pm 0.7\times 10^{8}$ ($\sim 8$% of the stellar mass). In addition to its double nucleus, the nuclear kinematics show pronounced asymmetries: a velocity-dispersion peak displaced by 6 pc from the galaxy center and a $\sim 16\,\mathrm{km\ s}^{-1}$ offset in the mean stellar line-of-sight velocity near the SMBH. We examine the origin of the 12 pc double nucleus and show that the observations favor an SMBH surrounded by an eccentric nuclear disk (END). One proposed formation pathway for ENDs involves a gravitational-wave (GW) recoil of the SMBH following a binary merger. Our orbit-superposition models contain $\sim 50\%$ retrograde stars at the edge of the nuclear region, in agreement with END-formation simulations. We infer a pre-merger mass ratio $q>0.15$ and a recoil kick of $\sim 340\,\mathrm{km\ s}^{-1}$. Our $N$-body simulations show that with such a kick the SMBH returns to the center within $\sim 30$ Myr. Its flat central core is also consistent with earlier “binary black hole scouring”. We test two alternative mechanisms—buoyancy-driven oscillations and a pre-merger SMBH binary—but neither reproduces the observed offsets, favoring the GW-kick scenario. Our $N$-body simulations also show that a prograde SMBH binary in a rotating host can become trapped in a corotation resonance, delaying coalescence. Although NGC 4486B is an old galaxy near the Virgo cluster center, its SMBH appears to have merged only recently, making its nucleus a rare nearby laboratory for post-merger SMBH dynamics.
@@
facilities: JWST(NIRSpec), HST(ACS/WFC), HST(WFPC2)
I Introduction
NGC 4486B (VCC1279) is a cE galaxy at the center of the Virgo Cluster, projected $7.3^{\prime}$ ($\sim 35$ kpc) from the massive elliptical M87, with a modest line-of-sight velocity difference of $\sim 200\,\mathrm{km}\,\mathrm{s^{-1}}$ that indicates a close dynamical association. NGC 4486B is characterized by a small effective radius of $2.33^{\prime\prime}$ ($\sim 0.19$ kpc) and an absolute magnitude of $M_{V}\sim 17.6$, with a total stellar mass of $_{*}\sim 6\times 10^{9}MM_{\odot}$ (Lauer et al., 1996; Ferrarese et al., 2006; Janz et al., 2016). Recently, we showed using JWST-NIRSpec IFU kinematical data and archival HST photometry, that NGC 4486B hosts a supermassive black hole (SMBH) with a mass of ${\cal M}_{BH}=(3.6\pm 0.7)\times 10^{8}\mbox{$M_{\odot}$}$, making this black hole significantly overmassive (${\cal M}_{BH}/M_{*}\sim 8$%) relative to expectations from black hole scaling relations (Tahmasebzadeh et al., 2025, hereafter Paper 1). This suggests that NGC 4486B is an extreme case of a tidally stripped galaxy, likely the compact core of a once significantly more massive system.
@@ I Introduction
Several formation pathways have been proposed for lopsided or double nuclear structures, including in-situ formation of eccentric disks through gas inflow and self-gravitating $m=1$ instabilities as well as perturbation-driven excitation of long-lived eccentric modes (e.g., Jacobs & Sellwood, 2001; Hopkins & Quataert, 2010). These mechanisms generally operate in systems with significant gas inflow or massive, self-gravitating nuclear disks. However, NGC 4486B is a gas-poor compact elliptical with an old stellar population and a central potential dominated by its SMBH, making such scenarios less well matched to its present-day properties.
@@
Lauer et al. (1996) investigated various explanations for the double nucleus in NGC 4486B. A dust absorption feature or an infalling globular cluster were ruled out due to the nearly identical morphologies in the $V$ and $I$ bands and the absence of significant color gradients across the nucleus. A bound stellar binary nucleus was likewise deemed unlikely, as its decay timescale due to dynamical friction would be short ($\lesssim 10^{7}$ yr), and there is no evidence for a recent galaxy–galaxy merger. Projection effects from unrelated foreground or background sources were also excluded based on the nearly identical colors and recession velocities of the two components (Kormendy et al., 1997).
One plausible explanation for the double nucleus in M31 and NGC 4486B is the presence of an eccentric nuclear disk (END) (Lauer et al., 1993, 1996), apsidally aligned stellar disk orbiting a central SMBH (Tremaine, 1995; Peiris & Tremaine, 2003; Bacon et al., 2001; Sambhus & Sridhar, 2002; Brown & Magorrian, 2013; Wernke & Madigan, 2021). In this model, the nearly Keplerian potential of the SMBH dominates the stellar dynamics, shaping a lopsided configuration where stars follow coherent, eccentric orbits. The apparent double nucleus arises from the asymmetric distribution of stellar density along these orbits: the brighter component corresponds to stars near their apocenters, where they move more slowly, while the fainter component is associated with stars near their pericenters, close to the black hole.
@@
A striking feature of NGC4486B is its flat central density core (see Section II.1). One of the best accepted mechanisms for producing such cores in early-type galaxies is via “binary-black-hole scouring” following the merger of two gas-poor galaxies, each with its own SMBH. The black holes (BHs) quickly sink to the common center of the galaxy via dynamical friction and become bound with separations of 10s of parsecs. In order to shrink the orbit further, the SMBH binary transfers its orbital kinetic energy to surrounding stars, via the $3$-body slingshot effect, preferentially ejecting stars with low angular momentum (e.g. Milosavljević & Merritt, 2003; Li et al., 2015) until the binary separation becomes small enough for gravitational wave (GW) radiation to dominate. High-resolution $N$-body simulations (Khan et al., 2012; Rantala et al., 2024) have shown that an SMBH binary can eject a stellar mass amounting to 252-5 times its total mass from the central region, thereby converting a central cusp of stars into a flatter central core (Rantala et al., 2018; Nasim et al., 2021; Khan et al., 2025).
@@
Numerical relativity simulations (e.g., Campanelli et al., 2007a, b) have shown that the anisotropic emission of GWs during the merger of two SMBH delivers a kick to the final black hole with a magnitude that depends on a variety of factors: the mass ratio $q=m_{2}/M_{1}$ of the secondary to the primary; whether the BHs have spin; the alignment of their spins to their mutual orbital plane; and whether the black hole spins are aligned or misaligned. While kicks as large as $\sim 2000-4000\,\mathrm{km\ s}^{-1}$ (Campanelli et al., 2007b) are possible, the actual distribution of kick velocities is uncertain and most kicks are probably $\lesssim 1000$$\mathrm{km\ s}^{-1}$(Gualandris & Merritt, 2008). Even a modest kick to the SMBH causes it to oscillate, transferring orbital energy to nearby stars, further flattening the central density profile (Gualandris & Merritt, 2008; Khonji et al., 2024).
@@
Recently, the END model has received renewed interest due to a suite of direct $N$-body simulations of nuclear disks surrounding SMBHs (Akiba & Madigan, 2021, 2023; Bright et al., 2024). These authors demonstrate that a GW recoil kick can naturally transform a circular stellar disk bound to an SMBH into an eccentric, apsidally aligned lopsided one. In this scenario, the stars that remain bound to the SMBH after the kick experience angular-momentum changes that place them on eccentric orbits, whose net alignment is oriented roughly orthogonal to the kick direction. For a sufficiently large kick, stars in the outer parts of the disk can even become retrograde in the rest frame of the SMBH.
@@
In this work, we investigate the photometric and kinematic signatures of the double nucleus in NGC 4486B and build upon our previous dynamical modeling results from Paper 1, as well as previous $N$-body simulations, to examine the origin and dynamical state of its END. We use previous analysis from Akiba & Madigan (2021, 2023); Akiba et al. (2025) and the observed properties of the double nucleus to estimate the kick velocity and likely mass-ratio of the black hole binary. We further investigate the dynamical evolution of the central SMBH using a self-consistent $N$-body realization constructed from the Schwarzschild orbit-superposition model in Paper 1. We simulate both the pre-merger evolution of an SMBH binary and the post-merger response of a recoiling SMBH within the host galaxy potential. These experiments probe the large-scale evolution of the SMBH in the global, axisymmetric potential inferred from the data and are not intended to model the internal structure or formation of the END. We simulate various recoil kicks to assess the timescale on which the black hole would sink to the center. In addition to GW recoil, we test other alternative mechanisms capable of displacing the SMBH and influencing the observed kinematics, such as buoyancy-driven displacements of the SMBH (Cole et al., 2012; Banik & van den Bosch, 2021, 2022; Dattathri et al., 2025b), and the pre-merger evolution of a prograde binary black hole merger (Holley-Bockelmann & Khan, 2015; Rasskazov & Merritt, 2016; Mirza et al., 2017; Khan et al., 2020). We also quantify the expected amplitude of SMBH Brownian motion (Merritt et al., 2007) to verify that the displacements seen in our simulations and inferred from observations exceed the stochastic wandering expected from finite stellar encounters Brownian motion.
@@
While nearly two decades of theoretical and simulation work have explored the consequences of binary black hole mergers, the resulting GW recoils, and SMBH Brownian motion in galaxies, NGC 4486B appears to be the first system exhibiting multiple observable signatures of a recent SMBH merger—making it a benchmark case for studying the dynamical aftermath of SMBH coalescence. Although NGC 4486B and M31 are among the best-studied examples, double nuclei are not unique. Debattista et al. (2006) examined a sample of dE galaxies from Lotz et al. (2004) imaged in the WFPC2 snapshot surveys GO-8600 and GO-6352 (PI: H. Ferguson) and found that VCC 1107, FCC 208, and VCC 128 show signs of double nuclei, with VCC 128 exhibiting two comparably bright components and the closest separation ($\sim 32$ pc). Tracing the formation and evolution of the eccentric disk in NGC 4486B and other galaxies may provide indirect constraints on the frequency of SMBH mergers – an essential parameter for interpreting the stochastic GW background measured by Pulsar Timing Arrays (PTAs) (Agazie et al., 2023; Ellis et al., 2024; EPTA Collaboration et al., 2023, 2024).
@@
The structure of this paper is as follows. In Section II, we present the HST and JWST observations of NGC 4486B and briefly summarize the Schwarzschild dynamical modeling from Paper 1. In Section III, we present the observational evidence for an END in NGC 4486B and compare the nuclear structure and kinematics of our Schwarzschild model with simulations of ENDs formed through GW recoil kicks. Using the observed properties of the double nucleus, we derive constraints on the black hole binary mass ratio and the required recoil velocity. In Section IV, we use direct $N$-body simulations to follow the post-kick trajectory of the SMBH and evaluate the timescale on which it returns to the galaxy center, as well as the kinematic signatures such a kick would imprint. Section V examines alternative mechanisms that could produce an off-center SMBH or similar kinematic asymmetries. Finally, in Section VI we summarize our findings and discuss their implications. Throughout this study, we adopt a distance of $16.3$ Mpc to NGC 4486B (Blakeslee et al., 2009).
II Observational Evidence and overview of Schwarzschild models in Paper 1
@@
In this section we briefly summarize the properties of photometric imaging from Hubble Space Telescope (HST) instruments WFPC2 and ACS/WFC and the spectroscopic data from JWST-NIRSpec IFU that were used to construct the Schwarzschild dynamical models of NGC 4486B presented in Paper 1 (Tahmasebzadeh et al., 2025). In addition, we briefly summarize the main results from that paper, including details about the best-fit Schwarzschild models that are used in this work.
II.1 Photometric data and image analysis
@@
In Paper 1, we used archival HST-ACS imaging in the F850LP filter (pixel scale $\sim 0.05\arcsec$ $\lesssim 4$ pc and PSF FWHM $\sim 0.9\arcsec$ $\lesssim 72$ pc) acquired as part of the ACS Virgo Cluster Survey (Côté et al., 2004). A model point spread function (PSF) was generated using the TinyTim software package (Krist et al., 2011) and used for all photometric analyses. The full photometric frame reaches out to $\sim$1 kpc in radius, with an effective radius of $2.33\arcsec$ (182 pc), extending well beyond the field of view of the NIRSpec-IFU observations. We also examine the WFPC2 F [^5] W image from Lauer et al. (1996), which provides slightly higher spatial resolution in the central region and reveals the double nucleus structure more distinctly.
Previous analyses of the morphology, isophotal parameters, and surface brightness profile of NGC 4486B showed that it exhibits a shallow surface brightness profile in its central region (Lauer et al., 1996; Faber et al., 1997; Ferrarese et al., 2006). Ferrarese et al. (2006) modeled its surface brightness using a single Sérsic component and obtained $n=2.73$ and $r_{e}=2.33\arcsec$. We repeated the 2D surface brightness fitting of the ACS F850LP image using GALFIT (Peng et al., 2010) models with one, two, three, and four Sérsic components. For a single component, the Sérsic index is $n=3.05$, slightly higher than in previous work. As the number of Sérsic components increases to three, the quality of the fit improved. Beyond this point, however, the model did not show substantial improvement (see Table 1). By adding more components, the Sérsic index of the inner region decreases to a range of n=1.21.4n=1.2-1.4. As might be expected due to the double nuclei, the inner region of NGC 4486B cannot be adequately described by single or multiple nested Sérsic components (even allowing for the components to have different centers). The major-axis position angles of all components across different models are similar and consistent with the position angle profile derived from the ellipse fitting to the isophotes. The ellipticity and position angle of the isophotes (outside the innermost $\sim 0.15\arcsec$ region occupied by the double nucleus) remain relatively constant with radius, indicating that the galaxy is close to axisymmetric overall. Similar results were obtained with the WFPC2 F [^5] W image.
@@
Figure 2: First row, right panel: Unsharp-masked WFPC2/F [^5] W image of the double nucleus in NGC 4486B, showing two peaks separated by $\sim$12 pc. First row, left and middle panels: Stellar kinematic maps of $v$ and $\sigma$ from JWST-NIRSpec IFU data, shown in Voronoi-binned cells. In all panels, surface-brightness contours from the deconvolved WFPC2/F [^5] W image are overplotted. The black dot marks the center of the faint peak (P2), which coincides with the $\sigma$ peak from NIRSpec and is identified as the likely location of the SMBH. The purple dot indicates the center of the bright peak (P1), coinciding with the center of the IFU data cube. The vertical dashed line marks the kinematic center. Second row: One-dimensional kinematic profiles of NGC 4486B derived from JWST-NIRSpec IFU data, showing the $|v|$, $\sigma$, and $|v|/\sigma$ values along the major axis intersecting both nuclei. Red points represent measurements and associated uncertainties from the right side of the kinematic maps, where $\sigma$ peaks. Blue points correspond to the left side of the galaxy, mirrored about the rotation axis for direct comparison. The gray shaded region marks the approximate radial extent of the eccentric disk inferred from photometric and kinematic features (corresponds to the white rectangle above).
In Paper 1, we used the multi-Gaussian expansion (MGE) method (Cappellari, 2002) to model the surface brightness of NGC 4486B. The details of the fit are provided in Table 1 of Paper 1. In the top panel of Figure 1, we present the projected luminosity density of NGC 4486B (black dots), along with the corresponding models: the MGE model (in red) and Sérsic models with one (green), two (blue), three (orange), and four (gray) components. The vertical dashed line indicates the PSF FWHM of the F850LP image. None of the models with multiple Sérsic components are able to replicate the flatness of the inner profile as well as the MGE model does in this region. To test whether the presence of the double nucleus affects the derived surface brightness profile, we allowed the centers of the Sérsic (or Gaussian) components to vary freely in our GALFIT modeling. In the best-fit model, the centers of all components are found to lie within $\leq 1$ pixel of each other. Likewise, in the standard MGE fitting, where all Gaussian components share a fixed center at the brightest peak, the result remains consistent. The choice of fixing or freeing the component centers therefore does not significantly affect the shape of the final luminosity profile. This is likely because, although the nuclear region is slightly lopsided rather than fully symmetric, the two peaks are separated by only $\simeq$ 3-4 pixels, comparable to the PSF FWHM. Consequently, the central light distribution appears smooth rather than distinctly double in the F850LP image, causing all model components to converge to nearly the same photocenter, which lies on the bright peak.
@@
As mentioned in Section I, the favored mechanism for core formation in elliptical galaxies is “core scouring by an SMBH binary”. The low-density core produced in this manner can extend up to five times the black hole binary influence radius (Rantala et al., 2018; Nasim et al., 2021; Khan et al., 2025). For the black hole in NGC 4486B ${\cal M}_{BH}=3.6\times 10^{8}$$M_{\odot}$ (Paper 1), the sphere of influence is estimated to be 203520-35 pc, which is comparable to the observed core radius.
The surface brightness fitting above ignored the double nucleus, which is barely resolved in the ACS/WFC F850LP image. In the higher spatial resolution of the HST WFC F [^5] W image (Lauer et al., 1996) however, it is clearly resolved (Figure 1 of Paper 1 gives both images). Figure 2 (first row, right panel) shows an unsharp-masked image of NGC 4486B, in which the double nucleus is revealed with striking clarity. The two peaks are distinctly separated by $\sim$12 pc, with the fainter component (P2) marked by a black dot while the brighter nucleus (P1) is marked by a purple dot. Photometrically, P1 has a central surface brightness of $\mu_{I}=12.89$ mag arcsec-2, while P2 has $\mu_{I}=13.00$ mag arcsec-2, with a minimum of $\mu_{I}=13.06$ mag arcsec-2 between them (Lauer et al., 1996). The white rectangle in this image encloses the nuclear region that we consider in greater detail below.
II.2 JWST NIRSpec-IFU kinematics